OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <8, 4>, v = <10, 7> pleasssseeeee :)

OpenStudy (jamesj):

How do you tell if two vectors are - parallel - orthogonal ?

OpenStudy (jamesj):

Well, two vectors are parallel if they are in the same direction or opposite directions. Hence just a scalar multiple of each other. That is, u and v are parallel if there is a number k such that u = kv Is there such a number k in this case?

OpenStudy (jamesj):

waiting for you to respond. I'm not going to say any more until you do

OpenStudy (anonymous):

sorry i was doing other parts of my test.. how do i know theres a number such as k?

OpenStudy (jamesj):

The question is this: IS there such a k? Well, if there were then u = kv <8, 4> = k<10, 7> i.e., 8 = 10k and 4 = 7k Can any ONE value of k satisfy both of those equations?

OpenStudy (anonymous):

no? lol i dont think so

OpenStudy (jamesj):

No indeed. Hence the vectors are not parallel. Now, what's the test for whether two vectors are orthogonal?

OpenStudy (anonymous):

i know that is if they're equal to zero...which theyre not so is it neither??

OpenStudy (jamesj):

No. Two vectors are orthogonal if their dot product is zero.

OpenStudy (anonymous):

right which i already tested and it isnt correct?

OpenStudy (jamesj):

Right. The dot product of u and v is u.v = 80 + 28 = 108 which is not zero. Hence the vectors u and v are not orthogonal.

OpenStudy (anonymous):

no neither! yay thanks a ton!! do you mind helping me with two more? lol

OpenStudy (jamesj):

Post new questions and if I have time I will. If they are questions about vectors being parallel or orthogonal, try and apply exactly the procedures we used here.

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, 8>, v = <-4, 8>

OpenStudy (jamesj):

By the properties of the dot product \[ u \cdot v = |u| \ |v| \ \cos \theta \] where \( \theta \) is the angle between u and v. So use this relationship and rearranging we have \[ \cos \theta = \frac{ u \cdot v } { |u| \ |v| } \] From this you can find the angle \( \theta \)

OpenStudy (anonymous):

okay so if that value is 68.. where do i go from there

OpenStudy (jamesj):

You have a text book or class notes yes? You definitely have worked examples of this sort of "angle between two vectors" problem. Have a look at that. You have found u.v, which is a start. Now you need to calculate the length of both vectors, |u| and |v|. And then calculate the value of u.v/(|u| |v|) Then you take the inverse cosine of that number.

OpenStudy (jamesj):

No, where did you get that? u = <-5, 8>, v = <-4, 8> Hence the length of u is \[ |u| = \sqrt{(-5)^2 + 8^2} = \sqrt{25 + 64} = \sqrt{89} \] Now do the same to find the length of v, |v|.

OpenStudy (anonymous):

oops that value is 84 not 68..

OpenStudy (anonymous):

are the lengths just u and v?

OpenStudy (jamesj):

By the way, u.v = 20 + 64 = 84 ... yes. I just showed you a length calculation. It is applying Pythagorus's theorem.

OpenStudy (anonymous):

\[\sqrt{80}\]

OpenStudy (jamesj):

Yes that's |v|. Now calculate u.v/(|u| |v|) because as we saw above, that is equal to the cosine of the angle between u and v.

OpenStudy (anonymous):

okay so 84/84.38

OpenStudy (anonymous):

ah got it! 5.4 degrees

OpenStudy (jamesj):

Sounds about right

OpenStudy (jamesj):

ok. I'm out of here. Good luck